From: Mon 11/10/08


  • UofC ATLAS meeting (3rd HEP) - skipped
  • FDR meeting (phone EsNet: call 1-510-883-7860, code 888726#; esnet video, ad hoc 888726)
  • troubleshooting AGLT2, directory creation error
  • compilation of MySQLdb
  • MWT2 meeting (Robert's office)
  • run Pathena job with 14.3.x or nightly (to test new Coll...GUID with Pathena)
  • leave early, 4pm


  • ATLAS ADCoS (prodsys shifters) meeting (phone CERN, Xavier Espinal) - skipped, overlapping
  • ATLAS TAG developers meeting (phone CERN, David Malon)
  • Take lead in closing AGLT2 LFC migration
    • new pilot from Paul fixing subdir problem
    • configuration problems (schedDB and proxies)
    • site back to online jobs running in the evening


  • ATLAS DB meeting (phone CERN) - skipped
  • ATLAS LFC meeting (phone Robert's office)
    • first prod job successfully finished at AGLT2, 40 in transferring state
    • ok to move to next sites
  • ATLAS facilities meeting (phone EsNet)
  • Help in MWT2_UC LFC migration (with Charles, Paul)
  • ATLAS shift


  • ATLAS shift
  • LFC migration at UC


  • LFC migration at UC
  • LFC problem troubleshooting at AGLT2

-- MarcoMambelli - 15 Oct 2008
Topic revision: r3 - 18 Oct 2008, MarcoMambelli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback